VIEWING DOCUMENTS AND IMAGES ESPECIALLY WITH EVINCE and XV

Carl Heiles January 17, 2016

Contents

You’ve created a wonderful written report as a pdf or ps file. Or you’ve created a beautiful figure as a pdf or ps file. Or you have a jpg image from your camera. How do you display and/or print these in Linux?

The most common display programs are in the following table. Our favorites are evince for documents and xv for images.

 pgm   img.ps img.eps img.pdf img.jpg  doc.pdf doc.ps  to     comments
                                 etc                  quit
evince    y      n       y       n        y       y  ctrl-w   no GUI menu
xv        y      y       y       y        y       y    q      pixels; some image editing
gv        y      y       y       n        y       y    q      gs + gui
display   y      n       n       y        n       n    q      good image editing
xpdf      n      n       y       n        y       n    q
gs        y      y       y       n        y       y    q      bare bones

Comments:

1. evince is the most capable for documents. The only problem is a temporary glitch: the GUI menu is unavailable. So you need to memorize a few basic keybindings, taken from evince keybindings.pdf The most important is to print use Ctrl p.

2. xv will handle almost any kind of file, especially image files. It has some image editing and file conversion; for a GUI, right-click. xv is not available on all Linus/UNIX installations. Disadvantage: xv pixellates the file before displaying it, so you can’t enlarge the image while retaining detail.

3. gv works on ps and pdf files only. It’s a little less nice than evince, but it has a nice GUI and is surely available on any Linux/UNIX platform.

4. display is meant for images. It has a GUI for lots of image editing options. It’s the modern replacement for xv.

5. xpdf is a basic, no-frills evince—but at least it has GUI, although it’s somewhat primitive,